Jack bought a pizza and drink for $2.65. If the pizza cost a penny more than 3 times the cost of drink, how much did each cost?

Answer by flame8855(424)

x + y = 6.25
x= 0.01 +y
2y + 0.01 = 6.25
2y = 6.24
y = 3.14
y = 3.11